# KVQuant: Towards 10 Million Context Length LLM Inference with KV Cache Quantization

### v3

@@ Appendix I Attention Sink-Aware Quantization Ablations
Table 17. Comparison of our NUQ datatype with and without sparsity against other data types on different models using the perplexity (PPL) measured on Wikitext-2. Non-uniform quantization (“nuq”) results are using pre-RoPE per-channel quantization for Keys. “gs64/128” refers to baseline experiments using grouping with group size 64/128. 2-bit results leverage per-matrix Q-Norm. Note that there is a slight variation in average bitwidth across models due to the differing hidden dimensions.
Method LLaMA-7b LLaMA-13b LLaMA-30b LLaMA-65b LLaMA-2-7b LLaMA-2-13b LLaMA-2-70b Mistral-7b Avg. Num. Bits
baseline 5.68 5.09 4.10 3.53 5.12 4.57 3.12 4.76 16
int4 5.98 5.32 4.34 3.73 5.66 5.01 3.31 4.97 4.00-4.02
int4-gs128 5.77 5.16 4.16 3.57 5.32 4.71 3.16 4.82 4.16
int4-gs64 5.73 5.14 4.14 3.56 5.25 4.66 3.14 4.80 4.31
nf4 5.87 5.23 4.25 3.63 5.47 4.90 3.22 4.91 4.00-4.03
nf4-gs128 5.77 5.17 4.17 3.58 5.30 4.71 3.16 4.83 4.25
nuq4 5.73 5.15 4.16 3.57 5.18 4.63 3.15 4.81 4.00-4.02
+ 0.1% outliers 5.71 5.12 4.15 3.55 5.16 4.61 3.14 4.80 4.04-4.06
+ 0.5% outliers 5.70 5.11 4.12 3.54 5.14 4.60 3.13 4.78 4.16-4.19
+ 1.0% outliers 5.70 5.11 4.12 3.54 5.14 4.59 3.13 4.78 4.32-4.35
int3 10.87 8.69 6.82 6.37 22.71 18.26 7.68 7.64 3.00-3.02
int3-gs128 6.17 5.47 4.44 3.78 6.15 5.34 3.33 5.16 3.15
int3-gs64 5.93 5.29 4.26 3.66 5.64 4.98 3.23 5.00 3.30
nf3 7.33 6.21 5.46 4.44 9.96 9.50 4.06 6.30 3.00-3.03
nf3-gs128 6.26 5.52 4.54 3.83 6.21 5.43 3.38 5.23 3.25
nuq3 6.01 5.34 4.41 3.74 5.49 4.83 3.26 5.03 3.00-3.02
+ 0.1% outliers 5.86 5.28 4.27 3.64 5.32 4.71 3.23 4.96 3.04-3.06
+ 0.5% outliers 5.79 5.15 4.19 3.60 5.22 4.66 3.17 4.86 3.16-3.19
+ 1.0% outliers 5.76 5.15 4.17 3.59 5.20 4.64 3.16 4.84 3.32-3.35
int2 11779 69965 1470 7272 4708 3943 976 573 2.00-2.02
int2-gs128 37.37 41.77 16.49 13.63 117.88 93.09 18.31 51.96 2.14
int2-gs64 11.09 9.84 6.60 5.54 25.69 26.83 5.93 12.47 2.28
nf2 3210.5 5785.6 2044.2 5367.3 13601 4035.6 3680.3 902.51 2.00-2.03
nf2-gs128 351.23 141.19 60.97 31.69 634.59 642.44 71.21 252.85 2.25
nuq2 8.17 7.29 7.05 27.17 9.75 29.25 4.98 7.33 2.00-2.02
+ 0.1% outliers 6.91 5.84 4.94 4.13 6.47 5.54 3.76 6.05 2.04-2.06
+ 0.5% outliers 6.26 5.52 4.48 3.86 5.67 5.05 3.35 5.33 2.16-2.19
+ 1.0% outliers 6.06 5.40 4.43 3.76 5.50 4.92 3.29 5.16 2.32-2.35
Table 18. Comparison of our NUQ datatype with and without sparsity against other data types on different models using the perplexity (PPL) measured on C4. Non-uniform quantization (“nuq”) results are using pre-RoPE per-channel quantization for Keys. “gs64/128” refers to baseline experiments using grouping with group size 64/128. 2-bit results leverage per-matrix Q-Norm. Note that there is a slight variation in average bitwidth across models due to the differing hidden dimensions.
Method LLaMA-7b LLaMA-13b LLaMA-30b LLaMA-65b LLaMA-2-7b LLaMA-2-13b LLaMA-2-70b Mistral-7b Avg. Num. Bits
baseline 7.08 6.61 5.98 5.62 6.63 6.05 4.97 5.71 16
int4 7.40 6.82 6.18 5.75 7.31 6.59 5.12 5.91 4.00-4.02
int4-gs128 7.16 6.67 6.02 5.65 6.87 6.20 5.00 5.76 4.16
int4-gs64 7.12 6.64 6.00 5.63 6.79 6.15 4.99 5.75 4.31
nf4 7.27 6.74 6.10 5.69 7.09 6.45 5.06 5.85 4.00-4.03
nf4-gs128 7.16 6.66 6.02 5.65 6.86 6.20 5.00 5.77 4.25
nuq4 7.13 6.65 6.02 5.64 6.70 6.11 5.00 5.75 4.00-4.02
+ 0.1% outliers 7.11 6.63 6.00 5.63 6.68 6.08 4.99 5.75 4.04-4.06
+ 0.5% outliers 7.10 6.62 5.99 5.62 6.66 6.07 4.98 5.73 4.16-4.19
+ 1.0% outliers 7.09 6.62 5.99 5.62 6.65 6.06 4.98 5.72 4.32-4.35
int3 12.97 10.95 9.13 8.27 30.14 28.57 16.00 8.84 3.00-3.02
int3-gs128 7.62 6.93 6.24 5.79 8.00 7.06 5.16 6.08 3.15
int3-gs64 7.34 6.78 6.11 5.70 7.29 6.59 5.08 5.92 3.30
nf3 8.90 7.84 7.43 6.37 14.92 13.75 5.96 7.27 3.00-3.03
nf3-gs128 7.65 6.99 6.29 5.82 8.03 7.12 5.24 6.16 3.25
nuq3 7.36 6.83 6.18 5.75 7.05 6.37 5.10 5.95 3.00-3.02
+ 0.1% outliers 7.24 6.72 6.08 5.68 6.87 6.20 5.06 5.88 3.04-3.06
+ 0.5% outliers 7.16 6.67 6.03 5.65 6.75 6.14 5.01 5.80 3.16-3.19
+ 1.0% outliers 7.14 6.66 6.02 5.64 6.72 6.11 5.00 5.77 3.32-3.35
int2 10892 100870 1411 7216 4708 4220 814 477 2.00-2.02
int2-gs128 43.49 56.25 21.07 17.05 113.49 97.04 23.67 50.73 2.14
int2-gs64 13.91 13.36 8.49 7.34 35.21 40.40 8.28 13.83 2.28
nf2 2850.1 4680.3 1617.1 5189.7 13081.2 4175.6 3216.9 1102.3 2.00-2.03
nf2-gs128 248.32 118.18 60.28 36.05 420.05 499.82 80.51 191.73 2.25
nuq2 10.28 9.05 9.27 60.87 15.16 43.77 7.24 8.40 2.00-2.02
+ 0.1% outliers 8.18 7.24 6.54 6.00 8.40 7.31 5.56 6.91 2.04-2.06
+ 0.5% outliers 7.50 6.91 6.21 5.78 7.28 6.53 5.16 6.23 2.16-2.19
+ 1.0% outliers 7.38 6.83 6.15 5.73 7.06 6.38 5.11 6.08 2.32-2.35
@@
Table 19. Full evaluation of our method for different models on all LLaMA, LLaMA2, and Mistral models using the perplexity on Wikitext-2. KVQuant results are using pre-RoPE per-channel quantization for Keys. Average bit widths assume a sequence length of 128K (ignoring context length limits for the models). Note that ATOM and FlexGen use 4-bit quantization with group sizes of 128 and 64 with uniform quantization, respectively, and we extend their methods to 3-bit and 2-bit quantization. We incorporate Q-Norm for 2-bit experiments, and we leverage Attention-Sink Aware quantization for all bit widths. We used post-RoPE quantization for all baseline methods since it achieves higher accuracy when quantizing Keys per-token as shown in Appendix N.
Method LLaMA-7b LLaMA-13b LLaMA-30b LLaMA-65b LLaMA-2-7b LLaMA-2-13b LLaMA-2-70b Mistral-7b Avg. Num. Bits
baseline 5.68 5.09 4.10 3.53 5.12 4.57 3.12 4.76 16
int4 5.98 5.32 4.34 3.73 5.66 5.01 3.31 4.97 4.00-4.02
nf4 5.87 5.23 4.25 3.63 5.47 4.90 3.22 4.91 4.00-4.03
ATOM-4bit 5.77 5.16 4.16 3.57 5.32 4.71 3.16 4.82 4.16
FlexGen-4bit 5.73 5.14 4.14 3.56 5.25 4.66 3.14 4.80 4.31
KVQuant-4bit 5.72 5.13 4.14 3.56 5.16 4.60 3.14 4.80 4.00-4.02
KVQuant-4bit-1% 5.69 5.10 4.11 3.54 5.13 4.59 3.13 4.78 4.32-4.35
int3 10.87 8.69 6.82 6.37 22.71 18.26 7.68 7.64 3.00-3.02
nf3 7.33 6.21 5.46 4.44 9.96 9.50 4.06 6.30 3.00-3.03
ATOM-3bit 6.17 5.47 4.44 3.78 6.15 5.34 3.33 5.16 3.15
FlexGen-3bit 5.93 5.29 4.26 3.66 5.64 4.98 3.23 5.00 3.30
KVQuant-3bit 5.89 5.25 4.30 3.63 5.35 4.71 3.18 4.99 3.00-3.02
KVQuant-3bit-1% 5.75 5.15 4.16 3.57 5.18 4.62 3.15 4.83 3.32-3.35
int2 11779 69965 1470 7272 4708 3943 976 573 2.00-2.02
nf2 3210.5 5785.6 2044.2 5367.3 13601 4035.6 3680.3 902.51 2.00-2.03
ATOM-2bit 37.37 41.77 16.49 13.63 117.88 93.09 18.31 51.96 2.14
FlexGen-2bit 11.09 9.84 6.60 5.54 25.69 26.83 5.93 12.47 2.28
KVQuant-2bit 7.15 5.81 4.84 4.38 6.95 11.63 3.52 7.05 2.00-2.02
KVQuant-2bit-1% 6.02 5.37 4.36 3.72 5.44 4.80 3.26 5.15 2.32-2.35
Table 20. Zero-shot MMLU evaluation results for LLaMA-7B and LLaMA-13B when incorporating 3-bit KV cache quantization. The value reported is the weighted accuracy across all tasks. We observe minimal accuracy degradation with 3-bit quantization.
Method LLaMA-7b LLaMA-13b Avg. Num. Bits
baseline 32.2% 36.6% 16
nuq3-1% 31.9% 36.2% 3.32-3.33
@@
Table 21 shows perplexity evaluation on Wikitext-2 for the LLaMA-7B model with uniform quantization, with Keys quantized pre-RoPE and post-RoPE. These results show that post-RoPE Key quantization is superior to pre-RoPE Key quantization when quantizing Keys per-token. This is because when rotating an outlier channel with large average magnitude and another channel with smaller average magnitude together, at some positions in the sequence part of the magnitude from the outlier channel will be shifted to the smaller channel. This partially mitigates the impact of the outlier channel on skewing the quantization range for some of the tokens in the sequence. As such, for our baseline comparisons, we use post-RoPE per-token Key quantization to serve as a stronger baseline.
@@
Figure 8 shows perplexity evaluation results with varying amounts of input context for 2-bit quantization. These results show that Q-Norm (and in particular, per-vector Q-Norm) can provide significant perplexity advantages for 2-bit quantization with long sequence length models. Additionally, for the LLaMA-2-70B-32K model, we observe perplexity degradation when going to a context length of 32K, likely due to error accumulation; however, this is largely mitigated by employing per-vector Q-Norm.
Figure 8. Perplexity results for the LLaMA-2-7B-32K model (uptrained for long sequence lengths using positional interpolation [3]) as well as the LLaMA-2-70B-32K LongLoRA model [4] on the Wikitext-2 dataset using 2-bit quantization, evaluated using different amounts of input context. Results are shown using Q-Norm (“q.2n”) as well as per-vector Q-Norm (“pv-qn”), and all results include 1% outliers. The memory reduction shown is for 2-bit quantization with per-vector Q-Norm. Our results show that Q-Norm is helpful for 2-bit quantization, and that per-vector Q-Norm provides additional benefits for 2-bit quantization with LLaMA-2-70B-32K. See Figure 4 for 4-bit and 3-bit quantization (which do not require Q-Norm).
Appendix P Mixed-Precision Quantization
An additional method to optimize compression performance is to consider mixed-precision quantization, where different layers are assigned different bit widths. This can allow for more accurate compression down to low bit widths due to differing sensitivities to quantization error of different layers. Finding the best bit precision distribution for different layers is intractable with brute force methods as the search space is exponentially large. We therefore aim to derive a metric to determine which layers can be quantized to reduced precision with minimal degradation in model accuracy, thereby enabling efficient one-shot mixed-precision bit assignments.
Prior work on mixed-precision quantization has leveraged the largest Hessian eigenvalue or the Hessian trace to assess which layers are most sensitive to quantization [9, 8]; however, due to the computational challenges of computing the full Hessian or even estimating Hessian eigenvalues, we instead leverage the Fisher information approximation for the Hessian (as derived in Appendix C). Using the diagonal Fisher information matrix along with the quantization error, we can use the following sensitivity metric for a given layer (with activation $A$ and quantized activation $A_{Q}$) to encompass both the quantization error as well as the Fisher information for that layer:
(6) $\Omega=\big{(}A-A_{Q}\big{)}^{\top}\mathcal{F}\big{(}A-A_{Q}\big{)}.$
We use the quantization error computed at the lower precision (as well as sensitivity information computed in fp16) to determine which layers were most sensitive to being quantized to the lower precision level. Figure 9 shows the mixed-precision perplexity results on the Wikitext-2 dataset for the LLaMA-7B and LLaMA-13B models. We show mixed-precision results using our sensitivity metric, and we include both quantization error-based mixed-precision assignment and the inverse of the selection order from our sensitivity metric as baselines. We see improved performance from incorporating sensitivity-based analysis for determining activation precisions when employing mixed-precision quantization. Our sensitivity metric therefore allows for efficiently determining an accurate mixed-precision assignment in order to trade off KV cache size and model accuracy.
Figure 9. Perplexity results for mixed-precision nuq4/3 quantization for the LLaMA-7B and LLaMA-13B models on the Wikitext-2 dataset, with model weights in fp16. Sensitivity-Weighted uses our metric for layer importance, Inverse Sensitivity-Weighted refers to selecting the lowest-importance layers using our metric, and Quantization Error uses the quantization error to sort layers without weighting using sensitivity information.
Appendix Q Kernel Implementation Details
@@
We implemented 4-bit lookup table-based kernels for matrix-vector multiplication between the Key or Value activations (packed as a lookup table (LUT) plus indices into the LUT per-element) and a full-precision activation vector. These kernels load the compressed Key and Value activations and dequantize them only as needed in order to minimize memory bandwidth utilization. All arithmetic is performed in fp16. The lookup table entries are the values of the sensitivity-weighted non-uniform datatype for that particular layer scaled according to the range of activations that need to be represented [6]. Note that it is possible to implement this using a single LUT shared across channels that is rescaled by a per-channel or per-token scaling factor and offset, but for simplicity we used a separate LUT per-channel or per-token for our initial implementation.
@@
Table 22 shows a detailed breakdown of kernel runtime, including how much time is spent packing vectors into the compressed format and how much time is spent on the dense and sparse matrix-vector multiplications. We find that even with 1% sparsity, we can attain significant speedups of up to 1.4$\times$ relative to the fp16 matrix-vector multiply kernels, demonstrating how our methodology facilitates efficient inference with a low-precision quantized KV cache.
Table 22. Average latency (in microseconds) for the Key and Value dense nuq4 kernels as well as for the sparse kernels (with 1% outliers), benchmarked on an A6000 GPU for the LLaMA-7B model. Benchmarking results are reported for different sequence lengths ($l$). fp16 matrix-vector multiplication latencies are included for reference, and the Key multiplication time also includes the time to apply RoPE to the newly appended Key vector. We find that our dense-and-sparse approach (even with 1% outliers) provides latency benefits relative to the fp16 baseline, even when accounting for the time to compress activations online.
@@
Key fp16 Matvec 66.4 116.1 402.3
Key (nuq4-1%) Dense Packing 2.4 2.4 2.4
Sparse Packing 4.5 4.5 4.5
Dense Matvec 42.8 75.5 283.5
Sparse Matvec 10.5 14.3 53.7
Total Latency 60.2 96.7 344.1
Value fp16 Matvec 56.0 104.2 389.3
Value (nuq4-1%) Dense Packing 2.0 2.0 2.0
Sparse Packing 4.2 4.2 4.2
Dense Matvec 26.3 63.1 226.3
Sparse Matvec 8.3 16.5 60.9
Total Latency 40.8 85.8 293.4

### v4

@@ Appendix I Attention Sink-Aware Quantization Ablations
Table 15. Comparison of our NUQ datatype with and without sparsity against other data types on different models using the perplexity (PPL) measured on Wikitext-2. Non-uniform quantization (“nuq”) results are using pre-RoPE per-channel quantization for Keys. “gs64/128” refers to baseline experiments using grouping with group size 64/128. Note that there is a slight variation in average bitwidth across models due to the differing hidden dimensions.
Method LLaMA Llama-2 Llama-3 Mistral-7B Avg. Num. Bits
7B 13B 30B 65B 7B 13B 70B 8B 70B
baseline 5.68 5.09 4.10 3.53 5.12 4.57 3.12 5.54 2.59 4.76 16
int4 5.98 5.32 4.34 3.73 5.66 5.01 3.31 7.89 14.03 4.97 4.00-4.02
int4-gs128 5.77 5.16 4.16 3.57 5.32 4.71 3.16 5.78 2.73 4.82 4.16
int4-gs64 5.73 5.14 4.14 3.56 5.25 4.66 3.14 5.68 2.66 4.80 4.31
nf4 5.87 5.23 4.25 3.63 5.47 4.90 3.22 6.42 3.86 4.91 4.00-4.03
nf4-gs128 5.77 5.17 4.17 3.58 5.30 4.71 3.16 5.84 2.75 4.83 4.25
nuq4 5.72 5.14 4.14 3.56 5.17 4.62 3.14 5.64 2.65 4.81 4.00-4.02
+ 0.1% outliers 5.70 5.12 4.12 3.55 5.15 4.60 3.13 5.63 2.66 4.79 4.04-4.06
+ 0.5% outliers 5.70 5.11 4.12 3.54 5.14 4.59 3.13 5.59 2.61 4.78 4.16-4.19
+ 1.0% outliers 5.69 5.10 4.11 3.54 5.13 4.59 3.13 5.57 2.60 4.78 4.32-4.35
int3 10.87 8.69 6.82 6.37 22.71 18.26 7.68 125 158 7.64 3.00-3.02
int3-gs128 6.17 5.47 4.44 3.78 6.15 5.34 3.33 7.50 4.11 5.16 3.15
int3-gs64 5.93 5.29 4.26 3.66 5.64 4.98 3.23 6.38 3.23 5.00 3.30
nf3 7.33 6.21 5.46 4.44 9.96 9.50 4.06 61.07 98.64 6.30 3.00-3.03
nf3-gs128 6.26 5.52 4.54 3.83 6.21 5.43 3.38 7.79 4.74 5.23 3.25
nuq3 5.94 5.32 4.34 3.68 5.39 4.82 3.23 6.10 2.99 4.98 3.00-3.02
+ 0.1% outliers 5.82 5.22 4.21 3.62 5.27 4.69 3.19 5.96 2.96 4.91 3.04-3.06
+ 0.5% outliers 5.76 5.16 4.16 3.59 5.20 4.65 3.16 5.74 2.69 4.84 3.16-3.19
+ 1.0% outliers 5.75 5.15 4.15 3.57 5.18 4.62 3.15 5.67 2.66 4.82 3.32-3.35
int2 11779 69965 1470 7272 4708 3943 976 2841 2164 573 2.00-2.02
int2-gs128 37.37 41.77 16.49 13.63 118 93.09 18.31 200 2092 51.96 2.14
int2-gs64 11.09 9.84 6.60 5.54 25.69 26.83 5.93 57.82 43.69 12.47 2.28
nf2 3210 5785 2044 5367 13601 4036 3680 30492 9486 903 2.00-2.03
nf2-gs128 351 141 60.97 31.69 635 642 71.21 1024 3091 253 2.25
nuq2 8.47 7.29 6.08 9.19 11.20 23.34 4.18 16.63 6.79 6.87 2.00-2.02
+ 0.1% outliers 6.82 5.72 4.83 4.00 6.38 5.33 3.54 7.97 4.77 5.83 2.04-2.06
+ 0.5% outliers 6.24 5.49 4.45 3.80 5.59 4.95 3.33 6.53 3.18 5.28 2.16-2.19
+ 1.0% outliers 6.05 5.39 4.41 3.72 5.47 4.85 3.28 6.29 2.95 5.14 2.32-2.35
Appendix M Full Perplexity Evaluation
Tables 16 and 17 show perplexity evaluation across all LLaMA, Llama-2, Llama-3, and Mistral models on Wikitext-2 and C4, respectively. These results demonstrate the benefits of our approach for KV cache compression across different models and bit widths.
@@
Table 16. Full evaluation of our method for different models on all LLaMA, Llama-2, Llama-3, and Mistral models using the perplexity on Wikitext-2. KVQuant results are using pre-RoPE per-channel quantization for Keys. Average bit widths assume a sequence length of 128K (ignoring context length limits for the models). Note that ATOM and FlexGen use 4-bit quantization with group sizes of 128 and 64 with uniform quantization, respectively, and we extend their methods to 3-bit and 2-bit quantization. We leverage Attention-Sink Aware quantization for all bit widths. We used post-RoPE quantization for all baseline methods since it achieves higher accuracy when quantizing Keys per-token as shown in Appendix N.
Method LLaMA Llama-2 Llama-3 Mistral-7B Avg. Num. Bits
7B 13B 30B 65B 7B 13B 70B 8B 70B
baseline 5.68 5.09 4.10 3.53 5.12 4.57 3.12 5.54 2.59 4.76 16
int4 5.98 5.32 4.34 3.73 5.66 5.01 3.31 7.89 14.03 4.97 4.00-4.02
nf4 5.87 5.23 4.25 3.63 5.47 4.90 3.22 6.42 3.86 4.91 4.00-4.03
ATOM-4bit 5.77 5.16 4.16 3.57 5.32 4.71 3.16 5.78 2.73 4.82 4.16
Flexgen-4bit 5.73 5.14 4.14 3.56 5.25 4.66 3.14 5.68 2.66 4.80 4.31
KVQuant-4bit 5.72 5.13 4.13 3.55 5.16 4.60 3.13 5.60 2.62 4.80 4.00-4.02
KVQuant-4bit-1% 5.69 5.10 4.11 3.54 5.13 4.58 3.13 5.56 2.60 4.78 4.32-4.35
int3 10.87 8.69 6.82 6.37 22.71 18.26 7.68 125 158 7.64 3.00-3.02
nf3 7.33 6.21 5.46 4.44 9.96 9.50 4.06 61.07 98.64 6.30 3.00-3.03
ATOM-3bit 6.17 5.47 4.44 3.78 6.15 5.34 3.33 7.50 4.11 5.16 3.15
FlexGen-3bit 5.93 5.29 4.26 3.66 5.64 4.98 3.23 6.38 3.23 5.00 3.30
KVQuant-3bit 5.87 5.25 4.25 3.63 5.34 4.71 3.18 5.84 2.72 4.98 3.00-3.02
KVQuant-3bit-1% 5.75 5.14 4.15 3.57 5.17 4.61 3.15 5.64 2.63 4.82 3.32-3.35
int2 11779 69965 1470 7272 4708 3943 976 2841 2164 573 2.00-2.02
nf2 3210 5786 2044 5367 13601 4036 3680 30492 9486 903 2.00-2.03
ATOM-2bit 37.37 41.77 16.49 13.63 118 93.09 18.31 200 2092 51.96 2.14
FlexGen-2bit 11.09 9.84 6.60 5.54 25.69 26.83 5.93 57.82 43.69 12.47 2.28
KVQuant-2bit 7.23 5.82 4.87 4.03 7.03 9.59 3.45 7.04 3.49 6.84 2.00-2.02
KVQuant-2bit-1% 6.01 5.36 4.35 3.70 5.41 4.78 3.26 6.04 2.82 5.14 2.32-2.35
Table 17. Full evaluation of our method for different models on all LLaMA, Llama-2, Llama-3, and Mistral models using the perplexity on C4. KVQuant results are using pre-RoPE per-channel quantization for Keys. Average bit widths assume a sequence length of 128K (ignoring context length limits for the models). Note that ATOM and FlexGen use 4-bit quantization with group sizes of 128 and 64 with uniform quantization, respectively, and we extend their methods to 3-bit and 2-bit quantization. We leverage Attention-Sink Aware quantization for all bit widths. We used post-RoPE quantization for all baseline methods since it achieves higher accuracy when quantizing Keys per-token as shown in Appendix N.
Method LLaMA Llama-2 Llama-3 Mistral-7B Avg. Num. Bits
7B 13B 30B 65B 7B 13B 70B 8B 70B
baseline 7.08 6.61 5.98 5.62 6.63 6.05 4.97 7.10 5.78 5.71 16
int4 7.40 6.82 6.18 5.75 7.31 6.59 5.12 8.79 14.36 5.91 4.00-4.02
nf4 7.27 6.74 6.10 5.69 7.09 6.45 5.06 7.93 6.58 5.85 4.00-4.03
ATOM-4bit 7.16 6.67 6.02 5.65 6.87 6.20 5.00 7.32 6.03 5.76 4.16
FlexGen-4bit 7.12 6.64 6.00 5.63 6.79 6.15 4.99 7.23 5.84 5.75 4.31
KVQuant-4bit 7.11 6.64 6.00 5.63 6.68 6.08 4.99 7.18 5.80 5.75 4.00-4.02
KVQuant-4bit-1% 7.09 6.62 5.99 5.62 6.64 6.06 4.98 7.12 5.78 5.72 4.32-4.35
int3 12.97 10.95 9.13 8.27 30.14 28.57 16.00 63.75 301 8.84 3.00-3.02
nf3 8.90 7.84 7.43 6.37 14.92 13.75 5.96 68.38 148 7.27 3.00-3.03
ATOM-3bit 7.62 6.93 6.24 5.79 8.00 7.06 5.16 9.00 6.72 6.08 3.15
FlexGen-3bit 7.34 6.78 6.11 5.70 7.29 6.59 5.08 7.92 6.18 5.92 3.30
KVQuant-3bit 7.25 6.72 6.06 5.68 6.89 6.18 5.03 7.43 5.90 5.92 3.00-3.02
KVQuant-3bit-1% 7.13 6.64 6.00 5.63 6.69 6.09 4.99 7.19 5.81 5.76 3.32-3.35
int2 10892 100870 1411 7216 4708 4220 814 2113 1977 477 2.00-2.02
nf2 2850 4680 1617 5190 13081 4176 3217 78331 7616 1102 2.00-2.03
ATOM-2bit 43.49 56.25 21.07 17.05 113 97.04 23.67 135 3734 50.73 2.14
FlexGen-2bit 13.91 13.36 8.49 7.34 35.21 40.40 8.28 50.78 42.27 13.83 2.28
KVQuant-2bit 8.52 7.32 6.67 5.96 9.49 15.36 5.30 10.06 6.45 7.63 2.00-2.02
KVQuant-2bit-1% 7.33 6.78 6.11 5.70 6.96 6.25 5.08 7.60 5.96 6.05 2.32-2.35
@@
Table 18 shows perplexity evaluation on Wikitext-2 for the LLaMA-7B model with uniform quantization, with Keys quantized pre-RoPE and post-RoPE. These results show that post-RoPE Key quantization is superior to pre-RoPE Key quantization when quantizing Keys per-token. This is because when rotating an outlier channel with large average magnitude and another channel with smaller average magnitude together, at some positions in the sequence, part of the magnitude from the outlier channel will be shifted to the smaller channel. This partially mitigates the impact of the outlier channel on skewing the quantization range for some of the tokens in the sequence. As such, for our baseline comparisons, we use post-RoPE per-token Key quantization to serve as a stronger baseline.
@@
Table 19 shows evaluation on Longbench [3] for the LLaMA-2-7B-32K model. The max input context length is set at 31500, and results using KIVI are also included for reference [25]. Note that we only ran LongBench evaluation using the 32K model since the majority of datasets have an average length greater than the context length of the non-extended models (i.e., 4K). The average input length was 12.2K tokens when using the LLaMA tokenizer (and truncating to fit within 32K context length), and hence most inputs would be heavily truncated for shorter context length models. Our results demonstrate that our method has little degradation relative to baseline fp16 inference, as well as improved accuracy for similar average bit width relative to KIVI.
Table 19. LongBench evaluation for the LLaMA-2-7B-32K model using KVQuant-3bit-1%. Comparisons with KIVI are included for reference, using the configuration with group size of 32 and 128-element fp16 residual [25]. Average bit widths are estimated for each approach assuming 12.2K context length, which was the average number of tokens across all tasks.
Task fp16 Baseline KIVI-2-gs32-r128 KVQuant-3bit-1%
Average Bit Width 16 3.17 3.33
Qasper 10.51 10.66 13.67
QMSum 16.99 16.36 16.53
MultiNews 22.15 23.37 16.54
TREC 71 71 71
TriviaQA 87.79 80.8 87.55
SAMSum 43.97 43.93 43.95
LCC 62.14 60.61 61.52
RepoBench-P 59.99 57.74 59.5
NarrativeQA 17.96 19.25 18.87
2WikiMQA 12.53 11.19 12.55
Musique 6.19 6.38 6.25
HotPotQA 12.55 12.48 12.07
Passage Retrieval 23 13.58 19.5
MultifieldQA-en 33.43 24.78 30.93
GovReport 29.65 27.05 27.1
Passage Count 1.50 1.50 1.75
Average Score 31.96 30.04 31.21
Appendix P Kernel Implementation Details
@@
We implemented 4-bit lookup table-based kernels for matrix-vector multiplication between the Key or Value activations (packed as a lookup table (LUT) plus indices into the LUT per-element) and a full-precision activation vector. These kernels load the compressed Key and Value activations and dequantize them only as needed in order to minimize memory bandwidth utilization. All arithmetic is performed in fp16. The lookup table entries are the values of the sensitivity-weighted non-uniform datatype for that particular layer scaled according to the range of activations that need to be represented [8].
@@
A second potential challenge is that the quantization dimension for the Keys and Values is not aligned with the reduction dimension for the respective matrix-vector operations. If we used per-channel or per-token lookup tables for dequantization, we would need to load from a different lookup table for each element as we iterate along the reduction dimension. However, since our implementation uses a single per-layer datatype that is rescaled per-vector, we can load a single scaling factor and zero-point for each channel / token and broadcast it across all threads (while doing a table lookup from the shared per-layer datatype, which is duplicated such that each thread can access it efficiently in parallel). The use of a single per-layer datatype that is rescaled per-vector therefore allows for much more efficient GPU kernel implementations in the context of per-channel Key / per-token Value quantization.
Table 20 shows a detailed breakdown of kernel runtime, including how much time is spent packing vectors into the compressed format and how much time is spent on the dense and sparse matrix-vector multiplications. For the fp16 baselines, we benchmarked runtime by averaging across 1000 iterations. When benchmarking our dense-and-sparse kernels (both for packing and matrix-vector multiplication), since the runtime of the sparse kernels can vary based on the sparsity pattern, we ran the LLaMA-2-7B-32K model and collected activations at each layer, and we then measured the average runtime across 1000 iterations for each layer in the model. We find that even with 1% sparsity, we can attain significant speedups of up to 1.7$\times$ relative to the fp16 matrix-vector multiply kernels, demonstrating how our methodology facilitates efficient inference with a low-precision quantized KV cache.
Table 20. Average latency (in microseconds) for the Key and Value dense nuq4 kernels as well as for the sparse kernels (with 1% outliers), benchmarked on an A6000 GPU for the LLaMA-2-7B-32K model. Benchmarking results are reported for different sequence lengths ($l$). fp16 matrix-vector multiplication latencies are included for reference, and the Key multiplication time also includes the time to apply RoPE to the newly appended Key vector. We find that our dense-and-sparse approach (even with 1% outliers) provides latency benefits relative to the fp16 baseline, even when accounting for the time to compress activations online.
@@
Key fp16 Matvec 33.3 59.1 219.4
Key (nuq4-1%) Packing 4.5 4.5 4.5
Dense Matvec 13.0 24.1 87.6
Sparse Matvec 8.1 11.2 34.2
Total Latency 25.6 39.9 126.3
Value fp16 Matvec 26.0 50.2 203.7
Value (nuq4-1%) Packing 4.1 4.1 4.1
Dense Matvec 10.0 17.8 62.0
Sparse Matvec 7.9 15.9 58.2
Total Latency 22.1 37.9 124.5